# Noise-induced synchronization in coupled quantum oscillators

### v2

@@ Abstract
We consider the quantum dynamics of a pair of coupled quantum oscillators coupled to a common correlated dissipative environment. The resulting equations of motion for both the operator moments and covariances can be integrated analytically using the Lyapunov equations. We find that for fully correlated and fully anti-correlated environments, the oscillators relax into a phase-synchronized state that persists for long-times when the two oscillators are nearly resonant and (essentially) forever if the two oscillators are in resonance. This can be traced to the symmetry of the Lindblad dissipator can lead to strong damping in one region of state space and underdamping in others. In the extreme cases of fully correlated or fully anti-correlated environments, specific regions of state space are fully decoupled from the environment. We also show that the environmental noise correlation leads to quantum entanglement, and all the correlations between the two oscillators are purely quantum mechanical in origin. This work provides a robust mathematical foundation for understanding how long-lived exciton coherences can be linked to vibronic correlation effects.
@@ I Introduction
While recovering from an illness in 1665, Christiaan Huygens observed that two identical pendulum clocks, when suspended from a common substantial beam, quickly synchronized with the same periods and amplitudes. [1, 2, 3, 4, 5, 6, 7, 8]. He described this intriguing behavior as ”an odd sympathy” and concluded that synchronization occurs due to the small vibrations transmitted through the support, which acts as a coupling mechanism between the pendula. Over time, these interactions facilitate an exchange of energy that synchronizes their motion. He presented these results to the Royal Society using a series of complex geometric arguments to explain the synchronization mechanism[8, 2].
The concept of synchronization has wide-ranging implications across various scientific fields, including the coordination of biological rhythms, the operation of power grids, and the development of synchronized communication systems [9, 10, 11, 12, 13, 14]. On the microscopic level, synchronization is crucial for in terms of developing quantum technologies that are robust against environmental effects such as decoherence, as well as in understanding the underlying mechanisms for the transport of quantum excitations in photosynthetic systems in which observed long-lived exciton coherences are linked to vibronic correlation effects. The enhanced excitonic coherence has a profound influence on the enhancement of exciton transport and delocalization [15, 16, 17, 18, 19, 20].
@@
We recently proposed that synchronization between pairs of qubits could be enhanced by coupling to a noisy ancillary environment with correlated or anti-correlated coupling.[21, 22] In these works, we show that noise correlation can help to preserve the fidelity and purity of a prepared state and determines the degree of phase synchronization such that anti-correlated noise leads to anti-phase synchronization and correlated noise leads to in-phase synchronization. In this work, we consider a pair of coupled quantum oscillators coupled to a common correlated dissipative environment. We show that correlations within the environment act as a shared resource that leads to entanglement in the steady-state solutions. The model is applicable to a wide range of physical systems from coherent exciton transport in light-harvesting complexes to quantum phase transitions in driven systems. Hence, we conclude that spontaneous synchronization is a universal hallmark of the interaction a quantum system with a correlated environment.
@@ II Lyapunov equations
in which $[a_{i},a_{j}^{\dagger}]=\delta_{ij}$ are bosonic operators with harmonic frequencies $\{\omega_{i}\}$ coupled via $g$ which allows the exchange of quanta from one local mode to the other. For the isolated system, the total number of excitations in the conserved since $[a_{1}^{\dagger}a_{1}+a_{2}^{\dagger}a_{2},(a_{1}^{\dagger}a_{2}+a_{2}^{\dagger}a_{1})]=0$. Further, we can remove the coupling term by transforming to a normal-mode representation without loss of generality. However, for purposes of our analysis we shall continue to work in a local-mode representation. The equations of motion take the general form
$\frac{d\rho}{dt}=-i[H,\rho]+D_{L}(\rho)$ (2)
where $D_{L}(\rho)$ is the Lindblad dissipator which has the general form
$D_{L}(\rho)=L\rho L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\rho\}$ (3)
where $L$ is a Lindblad operator describing the system/bath interactions.  [^1] We note that our analysis was done within Mathematica (v14.x) by defining a series bosonic algebra rules to generate and simplify the rather cumbersome expressions for multi-component boson systems. Included in this is a similar set of rules and transformations useful for multi-components spin systems. We include a link to a notebook with our derivations in the Data Availability section.
@@
in which $-1\leq\xi\leq 1$ serves as a correlation parameter. We use the eigenvalues and corresponding eigenvectors of this to construct new Lindblad operators corresponding to the correlated and anticorrelated components
@@
where $L_{1}$ is a Lindblad operator for oscillator 1 and $L_{2}$ is a Lindblad operator acting on oscillator 2. The total dissipator is then the sum of the symmetric and antisymmetric combinations. When $\xi=0$, the cross components vanish and we have uncorrelated oscillators. When $\xi=1$, the anti-symmetric contribution vanishes leaving only symmetric terms. This implies that symmetric normal modes will be more strongly damped by increasingly correlated environments, essentially leaving the anti-symmetric modes under damped to the extent that for fully correlated environments ($\xi\to 1$), symmetric normal modes of the system will be completely undamped. Likewise, in case of anti-correlated noise ($\xi=-1$), anti-symmetric normal modes are suppressed leading to in-phase synchronization.
In general, the correlation matrix $\Xi$ describes the topological adjacency between correlated Wiener processes such that $dW_{i}(t)dW_{j}(t^{\prime})=\Xi_{ij}\delta(t-t^{\prime})dt$ and may represent a connected graph belonging to a particular discrete symmetry group. Consequently, the eigenvectors of $\Xi$ can be classified according to the irreducible representations of that group.
Since our equations of motion generate a Gaussian map, we will instead work with the operator expectation values and covariances rather than with the density matrix itself. To establish our notation, we write the vector of moments of the system operators as ${\bf x}=\{\langle a_{1}\rangle,\langle a^{\dagger}_{1}\rangle,\langle a_{2}\rangle\langle a_{2}^{\dagger}\rangle\}$ and the covariances as
@@
It is straightforward to show that $\bf x$ and $\Theta$ evolve according to the following set of equations of motion
@@
where $W$ is the dynamical matrix and $f$ includes any source or sink terms. Similarly,
@@
where matrix $F$ depends only on the dissipative terms in the Liouville equation. These are referred to as the Lyapunov equations and are useful for understanding the stability and dynamics of chaotic systems[24].  [^2] Since these are equations of motion for the expectation values rather than the operators themselves, we use the adjoint dissipator $\langle\hat{O}D(\rho)\rangle={\rm tr}\left[{\hat{O}}\left(L\rho L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\rho\}\right)\right]=\langle L^{\dagger}\hat{O}L-\frac{1}{2}\{L^{\dagger}L,\hat{O}\}\rangle$ (11) in deriving these terms. For the coupled oscillators, we find the dynamical matrix is given by
$W=\left(\begin{array}[]{cccc}-i\omega_{1}&0&-ig-\xi\gamma_{12}/2&0\\ 0&+i\omega_{1}&0&ig-\xi\gamma_{12}/2\\ -ig-\xi\gamma_{12}/2&0&i\omega_{2}&0\\ 0&ig-\xi\gamma_{12}/2&0&+i\omega_{2}\end{array}\right)-\frac{\gamma}{2}I_{4}$ (16)
where $I_{4}$ is the $4\times 4$ identity matrix. Similarly,
$F=-\left(\begin{array}[]{cccc}\frac{\gamma}{2}+\gamma n_{1}&0&\xi\gamma\sqrt{n_{1}n_{2}}&0\\ 0&\frac{\gamma}{2}+\gamma n_{1}&0&\xi\gamma\sqrt{n_{1}n_{2}}\\ \xi\gamma\sqrt{n_{1}n_{2}}&0&\frac{\gamma}{2}+\gamma n_{2}&0\\ 0&\xi\gamma\sqrt{n_{1}n_{2}}&0&\frac{\gamma}{2}+\gamma n_{2}\\ \end{array}\right)$ (21)
@@
Starting with the moment equations
$\frac{\partial}{\partial t}\left(\begin{array}[]{c}\langle a_{1}\rangle\\ \langle a_{2}\rangle\end{array}\right)=-i\left(\begin{array}[]{cc}\omega_{1}-i\gamma/2&g-i\xi\gamma_{12}/2\\ g-i\xi\gamma_{12}/2&\omega_{2}-i\gamma/2\end{array}\right)\left(\begin{array}[]{c}\langle a_{1}\rangle\\ \langle a_{2}\rangle\end{array}\right)$ (28)
The non-Hermitian dynamical matrix has complex eigenvalues
$\lambda_{\pm}=\frac{1}{2}(\omega_{1}+\omega_{2}-i\gamma)\pm\frac{1}{2}\sqrt{(2g-i\xi\gamma_{12})^{2}+(\omega_{1}-\omega_{2})^{2}}$ (29)
where
$\gamma_{12}$ $=\gamma(\sqrt{(n_{1}+1)(n_{2}+1)}-\sqrt{n_{1}n_{2}})$ (30)
is a temperature dependent relaxation rate. If we take the case where the two oscillators are decoupled ($g=0$), synchronization becomes spontaneous when $\gamma_{12}\xi>(\omega_{1}-\omega_{2}$. Below the synchronization regime, both symmetric and anti-symmetric normal modes are equally damped at a rate given by $\gamma/2$. However, within the synchronization regime one normal mode becomes strongly damped while the other mode is more weakly damped. Only when the two modes are in resonance with $\omega_{1}=\omega_{2}$ does synchronization occur for all values of $\xi$. Similar regimes have been suggested by a series of recent papers concerning harmonic chains coupled by dissipative nearest neighbor terms [26, 27, 28, 29].
@@
Figure  1(a,b) shows the real and imaginary eigenvalues of the dynamical matrix in Eq. 28 for a $T=0K$ system with detuning $\delta=$ and relaxation rate $\lambda=$. Here, we set $\omega_{1}=1$ and $\omega_{2}=\omega_{1}+\delta$ to set the time scale of our system. The imaginary terms are also offset by $\lambda/2$. Below a critical value of $\xi$, the oscillators are damped as usual. We refer to this as the normal regime. At $\xi_{crit}=(\delta/\gamma_{12})$, we have an exceptional point where the two real normal mode frequencies coalesce, and one mode becomes less damped while the other becomes more strongly damped. As $\xi\to\pm 1$, one of the two normal modes will become completely decoupled from the dissipative environment.
@@
Figure 1: Real (a) and Imaginary (b) components of the eigenvalues of the non-Hermitian dynamical matrix in Eq. 28 for various parametric values: $\delta=\omega_{1}-\omega_{2}=0.2$, $\gamma=0.5$ and increasing values of exchange coupling $g=0$ to $g=0.04$ The system enters the spontaneous synchronized domain above the critical coupling $\xi_{crit}=\delta/\gamma$
@@
In Fig. 2 we show some representative trajectories for $\langle a_{1}(t)\rangle$ and $\langle a_{2}(t)\rangle$ for the case of zero detuning ($\delta=0$) and with various degrees of environmental correlations. We see clearly that anti-correlated noise ($\xi=-1$) leads to in-phase synchronization while fully correlated noise ($\xi=+1$) leads to anti-phase synchronization. Further more, at long time, the synchronized states are fully decoupled from the environment and do not decay as compared to the fully uncorrelated case which relaxes to zero on the time-scale dictated by $\gamma_{12}$.
@@
Figure 2: Representative dynamics for systems driven by anti-correlated (a), uncorrelated (b), and correlated (c) noise. (Parameters: $\delta=0$, $g=0.1$, $\gamma=0.1$)
We now turn our attention towards the equations of motion for the moments as given by Eq.10. The time-evolution is now entirely specified by the moments and covariance matrix. Representative trajectories for the moments are given in Fig.1(a-c). Here, we initiate the dynamics from the same initial condition (corresponding to to two displaced oscillators), however, with differing degrees of environmental correlation. In Fig. 1(a) and (c), phase synchronization or anti-synchronization is achieved after only a few oscillation periods, after which the dynamics settles into a steady-state limit cycle. In contrast, for uncorrelated environments, the two essentially independent oscillators relax as one would expect.

### v3

@@ Abstract
We consider the quantum dynamics of a pair of coupled quantum oscillators coupled to a common correlated dissipative environment. The resulting equations of motion for both the operator moments and covariances can be integrated analytically using the Lyapunov equations. We find that for fully correlated and fully anti-correlated environments, the oscillators relax into a phase-synchronized state that persists for long-times when the two oscillators are nearly resonant and (essentially) forever if the two oscillators are in resonance. We identify an exceptional point that indicates the onset of broken symmetry between an unsynchronized and synchronized dynamical phase of the system as correlations within the environment are increased. We also show that the environmental noise correlation leads to quantum entanglement, and all the correlations between the two oscillators are purely quantum mechanical in origin. This work provides a robust mathematical foundation for understanding how long-lived exciton coherences can be linked to vibronic correlation effects.
@@ I Introduction
The concept of synchronization has wide-ranging implications across various scientific fields, including the coordination of biological rhythms, the operation of power grids, and the development of synchronized communication systems [9, 10, 11, 12, 13, 14]. On the microscopic level, synchronization is crucial for developing quantum technologies that are robust against environmental effects such as decoherence [15], as well as in understanding the underlying mechanisms for the transport of quantum excitations in photosynthetic systems in which observed long-lived exciton coherences are linked to vibronic correlation effects. The enhanced excitonic coherence has a profound influence on the enhancement of exciton transport and delocalization [16, 17, 18, 19, 20, 21, 22, 23].
We recently proposed that synchronization between pairs of qubits could be enhanced by coupling to a noisy ancillary environment with correlated or anti-correlated coupling.[24, 25] In these works, we show that noise correlation can help to preserve the fidelity and purity of a prepared state and determines the degree of phase synchronization such that anti-correlated noise leads to anti-phase synchronization and correlated noise leads to in-phase synchronization. In this work, we consider a pair of coupled quantum oscillators coupled to a common correlated dissipative environment. We show that correlations within the environment act as a shared resource that leads to entanglement in the steady-state solutions. The model is applicable to a wide range of physical systems from coherent exciton transport in light-harvesting complexes to quantum phase transitions in driven systems. Hence, we conclude that spontaneous synchronization is a universal hallmark of the interaction between a quantum system and a correlated environment.
@@ II Lyapunov equations
in which $[a_{i},a_{j}^{\dagger}]=\delta_{ij}$ are bosonic operators with harmonic frequencies $\{\omega_{i}\}$ coupled via $g$ which allows the exchange of quanta from one local mode to the other. For the isolated system, the total number of excitations is conserved since $[a_{1}^{\dagger}a_{1}+a_{2}^{\dagger}a_{2},(a_{1}^{\dagger}a_{2}+a_{2}^{\dagger}a_{1})]=0$. Further, we can remove the coupling term by transforming to a normal-mode representation without loss of generality. However, for purposes of our analysis we shall continue to work in a local-mode representation. The equations of motion take the general form
$\frac{d\rho}{dt}=-i[H,\rho]+\sum_{i}D_{i}(\rho)$ (2)
where $D_{i}(\rho)$ is the total Lindblad dissipator for oscillator $i$ which has the general form
$D_{i}(\rho)=L_{i}\rho L_{i}^{\dagger}-\frac{1}{2}\{L_{i}^{\dagger}L_{i},\rho\}+L^{\prime}_{i}\rho{L^{\prime}}_{i}^{\dagger}-\frac{1}{2}\{{L^{\prime}}_{i}^{\dagger}L^{\prime}_{i},\rho\}$ (3)
where $L_{i}$ and $L^{\prime}_{i}$ are Lindblad operators describing the local system/bath interactions associated with oscillator $i$ [^1] We note that our analysis was done within Mathematica (v14.x) by defining a series bosonic algebra rules to generate and simplify the rather cumbersome expressions for multi-component boson systems. Included in this is a similar set of rules and transformations useful for multi-components spin systems. We include a link to a notebook with our derivations in the Data Availability section. For these, we use
$L_{i}$ $=\sqrt{\gamma(\overline{n}_{i}+1)}a_{i}$ (4)
$L^{\prime}_{i}$ $=\sqrt{\gamma\overline{n}_{i}}a^{\dagger}_{i}$ (5)
for the local Lindblad terms where $\gamma$ give the thermalization rate and $\overline{n}_{i}$ is the thermal population.
@@
in which $-1\leq\xi\leq 1$ serves as a correlation parameter. In general, the correlation matrix $\Xi$ describes the graphical adjacency between Wiener processes such that $dW_{i}(t)dW_{j}(t^{\prime})=\Xi_{ij}\delta(t-t^{\prime})dt$.
We use the eigenvalues and corresponding eigenvectors of $\Xi$ to construct new Lindblad operators
$L_{\alpha}=\sqrt{\xi_{\alpha}}\sum_{i}T_{\alpha i}L_{i}$ (9)
where $T^{\dagger}\Xi T={\rm diag}\{\xi_{\alpha}\}$. In general, the correlation matrix, $\Xi$, describes the topological adjacency between correlated Wiener processes such that $dW_{i}(t)dW_{j}(t^{\prime})=\Xi_{ij}\delta(t-t^{\prime})dt$ and may represent a connected graph belonging to a particular discrete symmetry group. Consequently, the eigenvectors of $\Xi$ can be classified according to the irreducible representations of that group. For the special case of a $k$-regular graph (that is a graph with each node having $k$ edges and $\Xi=I+\xi A$, where $A$ is the adjacency matrix for an undirected graph and $I$ is the identity matrix), one can always find one eigenstate of $\Xi$ with an eigenvalue equal to 0. For example, for a cyclic 2-regular graph with $N$ nodes
$\xi_{\alpha}=1+2\xi\cos(\frac{2\pi}{N}\alpha)$ (10)
where $\alpha=0,1,\lfloor N/2\rfloor)$. The eigenvectors with $11\leq\alpha\leq\lceil N/2\rceil-$ 1 are doubly degenerate. For even $N$, the $\alpha=0$ and $\alpha=N/2$ eigenvalues are 00 and $2\xi$, respectively. with the corresponding eigenvectors are $(1,1\cdots,1)$ and $(1,-1,1,-1\cdots)$. However, if $\xi=1$, the $\alpha=0$ eigenvector is the totally anti-symmetric eigenvector and if $\xi=-1$, then the $\alpha=0$ eigenvector is totally symmetric. Consequently, the symmetry of the 00-mode implies that corresponding irreducible representations of the system will be totally decoupled from the environment.
In our calculations which follow, we use the following set of local Lindblad operators:
$L_{i}=\sqrt{\gamma(\overline{n}_{i}+1)}a_{i}$ (11)
$L_{i}^{\prime}=\sqrt{\gamma\overline{n}_{i}}a^{\dagger}_{i}$ (12)
where $L_{i}$ describes the relaxation and $L^{\prime}_{i}$ describes the thermal excitations. The relaxation rate is specified by $\gamma$ and $\overline{n}_{i}$ is the equilibrium thermal population of oscillator $i$. We use the eigenvalues and corresponding eigenvectors of $\Xi$ to construct new Lindblad operators corresponding to the correlated and anticorrelated components
@@
$L^{\prime}_{\pm}=\sqrt{\frac{1\pm\xi}{2}}(L^{\prime}_{1}\pm L^{\prime}_{2}).$ (14)
The total dissipator now includes cross terms involving $L_{1}$ and $L_{2}$ and similarly $L^{\prime}_{1}$ and $L^{\prime}_{2}$ arising from correlation between the two local environments. When $\xi=0$, the cross components vanish and dissipation channels are independent of each other. When $\xi=1$, the anti-symmetric contribution vanishes leaving only symmetric terms $L_{+}$ and $L^{\prime}_{+}$. Likewise, in case of anti-correlated noise ($\xi=-1$), symmetric terms vanish leaving only the anti-symmetric terms $L_{-}$ and $L^{\prime}_{-}$.
In principle, one can also write $H$ in terms of the normal modes of the oscillators themselves. For the the two oscillator system we have a symmetric and an anti-symmetric normal mode. Consequently when $\xi>0$ (more correlated environment), the symmetric normal mode will be strongly damped compared to the anti-symmetric mode such that in the extreme case of $\xi<0$ the anti-symmetric mode is completely undamped and vice-versa for $\xi=0$.
Since our equations of motion generate a Gaussian map, we will instead work with the operator expectation values and covariances rather than with the density matrix itself. To establish our notation, we write the vector of moments of the system operators as ${\bf x}=\{\langle a_{1}\rangle,\langle a^{\dagger}_{1}\rangle,\langle a_{2}\rangle,\langle a_{2}^{\dagger}\rangle\}$ and the covariances as
@@
$\Theta+i\Omega\geq 0$ (16)
where $\Omega$ is the symplectic form
The covariance matrix and moments can be used to construct the Wigner function in multi-variable form
${\cal{W}}(\alpha)=\frac{1}{(2\pi)^{N}\sqrt{|\Theta|}}\exp\left[-\frac{1}{2}(\alpha-{\bf x})^{\dagger}\cdot\Theta^{-1}\cdot(\alpha-{\bf x}).\right]$ (20)
where $N$ is the number of oscillators and $\alpha$ is a continuous multi-dimensional complex variable whose real and imaginary components are related to the position and momentum conjugate variables. Note, that the Husimi-Q function and Glauber-Sudarshan P functions are similarly constructed; however, the Q-function has a covariance $\Theta_{Q}=\Theta+I$ while the P-function uses $\Theta_{P}=\Theta-I$.
The time-evolution of any observable, $\hat{O}$, is given by the master equation
$\frac{d\langle\hat{O}\rangle}{dt}=i\langle[H,\hat{O}]\rangle+\langle\overline{D}_{L}(\hat{O})\rangle$ (21)
Since these are equations of motion for the expectation values rather than the operators themselves, we use the adjoint dissipator
$\overline{D}_{L}(\hat{O})=\langle\hat{O}D_{L}(\rho)\rangle$ $={\rm tr}\left[{\hat{O}}\left(L\rho L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\rho\}\right)\right]$
$=\frac{1}{2}\langle L[\hat{O},L^{\dagger}]+[L,\hat{O}]L^{\dagger}\rangle$ (22)
in deriving these terms.
It is straightforward to show that the moments, $\bf x$, and covariance, $\Theta$, evolve according to the following set of equations of motion
@@
where $W$ is the dynamical matrix and $f$ includes any source or sink terms. Similarly the covariances evolve according to
@@
where matrix $F$ depends only on the dissipative terms in the Liouville equation. These are referred to as the Lyapunov equations and are useful for understanding the stability and dynamics of chaotic systems[29].
For the coupled oscillators, we find the dynamical matrix is given by
$W=\left(\begin{array}[]{cccc}-i\omega_{1}&0&-ig-\xi\gamma_{12}/2&0\\ 0&+i\omega_{1}&0&ig-\xi\gamma_{12}/2\\ -ig-\xi\gamma_{12}/2&0&-i\omega_{2}&0\\ 0&ig-\xi\gamma_{12}/2&0&+i\omega_{2}\end{array}\right)-\frac{\gamma}{2}I_{4}$ (29)
where $I_{4}$ is the $4\times 4$ identity matrix and
$\gamma_{12}$ $=\gamma(\sqrt{(\bar{n}_{1}+1)(\bar{n}_{2}+1)}-\sqrt{\bar{n}_{1}\bar{n}_{2}})$ (30)
is a temperature dependent relaxation rate. Similarly,
$F=-\left(\begin{array}[]{cccc}\frac{\gamma}{2}+\gamma\overline{n}_{1}&0&\xi\gamma\sqrt{\overline{n}_{1}\overline{n}_{2}}&0\\ 0&\frac{\gamma}{2}+\gamma\overline{n}_{1}&0&\xi\gamma\sqrt{\overline{n}_{1}\overline{n}_{2}}\\ \xi\gamma\sqrt{\overline{n}_{1}\overline{n}_{2}}&0&\frac{\gamma}{2}+\gamma\overline{n}_{2}&0\\ 0&\xi\gamma\sqrt{\overline{n}_{1}\overline{n}_{2}}&0&\frac{\gamma}{2}+\gamma\overline{n}_{2}\\ \end{array}\right)$ (35)
The derivation of these equations is straightforward and was facilitated by a Mathematica package we developed for doing multi-boson operator algebra. This package is publically available from the URL given in the Data Availability Statement below.
@@
The dynamical matrix, $W$, is non-Hermitian and has complex eigenvalues
$\lambda_{\pm}=\left(\frac{i}{2}(\omega_{1}+\omega_{2})-\gamma\right)\pm\frac{i}{2}\sqrt{(2g-i\xi\gamma_{12})^{2}+(\omega_{1}-\omega_{2})^{2}}$ (36)
If we take the case where the two oscillators are decoupled ($g=0$), synchronization becomes spontaneous when $\xi>(\omega_{1}-\omega_{2})/\gamma_{12}$.
This defines the critical value of $\xi$ necessary for synchronization to occur. As $\xi\to\pm 1$, one of the two normal modes will become completely decoupled from the dissipative environment, leading to the formation of undamped oscllations in which the two oscillators are completely in-phase or out-of-phase with each other. This is the synchronized (or antisynchronized) regimes.
@@
Figure  1(a,b) shows the real and imaginary eigenvalues of the dynamical matrix for a system at $T=0K$ with detuning $\delta=0.2$ and relaxation rate $\gamma=0.5$. Here, we set $\omega_{1}=1$ and $\omega_{2}=\omega_{1}+\delta$ to set the time scale of our system. Both the real and imaginary components of the eigenvalues are shifted as indicated on the plots. Here, the eigenvalues of $W$ appear to exhibit an exceptional point at $\xi_{crit}=(\delta/\gamma_{12})$, where the two frequencies coalesce ($\text{Im}(\lambda_{+})=\text{Im}(\lambda_{-}$)) and one mode becomes less damped while the other becomes more strongly damped. An exceptional point signals a breaking of parity-time (${\cal P}{\cal T}$) symmetry. [30, 31, 32, 33, 34, 35, 36] In the normal regime, the system exhibits a balanced distribution of gain and loss. When this balance is disrupted, the system can transition into a phase where the ${\cal P}{\cal T}$ symmetry is broken. In a ${\cal P}{\cal T}$-symmetry broken system, the fluctuations can become asymmetric, meaning the probability of a fluctuation in one direction can be significantly different from the probability of a fluctuation in the opposite direction, leading to a violation of the fluctuation dissipation theorem.[37]
@@
Figure 1: Real (a) and Imaginary (b) components of the eigenvalues of the non-Hermitian dynamical matrix in Eq. 29 for various parametric values: $\delta=\omega_{1}-\omega_{2}=0.2$, $\gamma=0.5$ and increasing values of exchange coupling $g=0$ to $g=0.04$ The system enters the spontaneous synchronized domain above the critical coupling $\xi_{crit}=\delta/\gamma$. (c) Analytical continuation of (a) onto the complex $\xi$-plane for showing the presence of a branch cut starting at ±$\pm i\xi_{crit}$. (d) Slice through (c) along $|\xi_{crit}|$ passing through the exceptional points at $\phi=\pm\pi/2$ for $g>0$. Solid: $\text{Im}(\lambda_{+})$. Dashed: $\text{Im}(\lambda_{-})$.
Setting $g\neq 0$ lifts the degeneracy between the two oscillators and there is no longer an exceptional point along the real $\xi$-axis. However, one can find the exceptional point by performing an analytical continuation of $\xi\to|\xi|e^{i\phi}$ in $W$. This places the exceptional point along the imaginary $\xi$-axis ($\phi=\pi/2$) for all values of $g$. Further, this is the start of a seam where $\text{Im}(\lambda_{+})=\text{Im}(\lambda_{-})$ along the imaginary axis staring at ±$\pm i\xi_{crit}$. This is illustrated in Fig. 1(c) where we show the analytical continuation of $\text{Im}(\lambda_{\pm})$ on the complex $\xi$-plane. Consequently, one can adiabatically transition between synchronized and anti-synchronized regimes by transforming the system along $|\xi|e^{i\phi}$. While this may not be feasible in a material system, it may be possible to construct light-matter or fully optical systems that exhibit this behavior.[38] Lastly, we note that $g<0$ swaps the two regimes such that $\xi<0$ leads to anti-synchronized and $\xi>0$ produces synchronized dynamics.
@@
In Fig. 2 we show two representative trajectories for $\langle a_{1}(t)\rangle$ and $\langle a_{2}(t)\rangle$ for the case of zero detuning ($\delta=0$) and with various degrees of environmental correlations. Here, we simply chose random initial conditions for the $\langle a_{1}\rangle$ and $\langle a_{2}\rangle$ moments. We see clearly that anti-correlated noise ($\xi=-1$) leads to in-phase synchronization while fully correlated noise ($\xi=+1$) leads to anti-phase synchronization. Further more, at long time, the synchronized states are fully decoupled from the environment and do not decay as compared to the fully uncorrelated case which relaxes to zero on the time-scale dictated by $\gamma_{12}$. Similar regimes have been suggested by a series of recent papers concerning harmonic chains coupled by dissipative nearest neighbor terms [39, 40, 41, 42].
Figure 2: Representative moment trajectories for pairs of equivalent
@@
oscillators driven by (a)anti-correlated $(\xi=-1)$, (b) uncorrelated $(\xi=0)$, and (c) correlated ($\xi=+1)$ noise. (Parameters: $\delta=0$, $g=0.1$, $\gamma=0.1$.)
We can quantify phase synchronization by examining the phase-locking value (PLV), taking as the average phase difference between $\langle a_{1}(t)\rangle$ and $\langle a_{2}(t)\rangle$,
[…]
where $\phi_{i}(t)=\arg(\langle a_{i}(t)\rangle)$ is the phase-angle associated with the complex-valued moment. For this, we first integrate the equation of motion for some sufficiently long time, $t_{1}\gg 1/\gamma$. We then sampled over 5000 initial phase differences, varying $g$ to accumulate the data shown in Fig. 3(a). For $g=0$, synchronization does not occur for $-\xi_{crit}\leq\xi\leq\xi_{crit}$. Outside this regime, the PLV goes to $+1$ for $-\xi_{crit}\geq-1$ and to $-1$ for $\xi_{crit}\leq+1$, indicating the presence of synchronized or anti-synchronized states. For $g\neq 0$, phase synchronization or anti-synchronization readily occurs, except when $\xi=0$. For $-\xi_{crit}\leq\xi\leq+\xi_{crit}$ and $g=0$, the two oscillators are asymptotically unsynchronized. The oscilations in the computed PLV values are due to numerical sampling of the initial conditions and finite time integration of the PLV integral. These can be attributed to the competition between gain and loss mechanisms while the system is in the non-synchronized regime.
When the coupling $g\neq 0$, the PLVs show obvious dips close to the critical value of the correlation parameter indicating the “competition” between relaxation and synchronization. In Fig 3(b) we set $\delta=0$ and vary the relaxation rate $\gamma$ to illustrate how the switch between synchronized to anti-synchronized behavior depends on the coupling to the environment, again averaging over initial conditions of the moments. Weakening the system/bath coupling $\gamma$ weakens the synchronization effects.
Figure 3: (a) Average phase-locking value (PLV) versus $\xi$ for $\delta=0.2$ and off-diagonal coupling $0\leq g\leq 0.05$. (b)Average PLV for $\delta=0$ and $g=0$, varying the relaxation constant $\gamma$.